OpenStudy (anonymous):

How many full cycles of the function y = 4 cos 0.5x appear in 4(pi) radians?

OpenStudy (anonymous):

1?

geerky42 (geerky42):

Period is \(0.5x = 2\pi = \dfrac{1}{2}x \Rightarrow x =\) \(4\pi\) So it appears one time for every \(4\pi\).